OpenStudy (anonymous):

Solve for x. [3x+5]/[(x-1)(x^4+7)]=0

OpenStudy (anonymous):

-5/3

OpenStudy (anonymous):

can you show the steps?

OpenStudy (anonymous):

clear denominators

OpenStudy (anonymous):

that will give 3x+5=0 can you solve that?

OpenStudy (anonymous):

no idk how

OpenStudy (anonymous):

3x+5=0 subtract 5 from both sides 3x=-5 divide both sides by 3 x=-5/3

OpenStudy (anonymous):

what about the domain restrictions?

OpenStudy (anonymous):

They wouldn't have a play here. -5/3 is absolutely right.

OpenStudy (anonymous):

(x-1)(x^4+7) cant be = 0

OpenStudy (anonymous):

Seems a bit to advanced I'd assume D is real numbers.

OpenStudy (anonymous):

lol ya